Bank ABC is advertising its current rate for car loans. The sign in the window says 4% APR with monthly compounding. Part A) If instead the bank advertised the effective annual rate (EAR), what would the sigrn say? Part B) What is the effective monthly interest rate? Part C) What is the effective quarterly interest rate? Part D) What is the effective interest rate per decade?
Answer #1

a) Effective annual rate = (1+4%/12)12 -1 = 4.07%

b) Effective monthly rate = 4%/12 = 0.33%

c) effective Quarterly rate = (1+0.3%)3 -1 = 1.003%

d) Effective interest rate per decade = (1+4%/12)12*10 -1 = 49.08%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well